AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri Apr 4 17:16:56 2025
Author
*
:
Subject
*
:
> 11.00 No WLAN/WiFi > came-gsi private LAN appears to be OK > > AIDA FEE64 power ON > > 11.09 FEE64 temps - attachment 1 > aida06 low, otherwise OK > > 11:10 CARME IE514 4.0e-11 mbar, temp unknown > IE514 OFF > > 11.22 Detector bias ON - attachment 2 > found ch 2 up ramp set to 63V/s! changed to 1V/s > > 11.30 system wide checks OK except > aida02 & aida06 fail WR decoder status > > WR ts control - attachments 3-4 > > > 12.02 ASIC settings updated for aida09-aida16 > ASIC settings DB key 2024Jan20-12:02:08 > > Unable to save DB key to Options - use ASIC settings restore as workaround > > 12.30 Layouts for 16x FEE64s created - directory ~/LayOut/16xFEE64 > Note layouts include spectra from aida07, aida08 and aida16 so an error message will be generated but it can be ignored > > > 12.40 per FEE64 Rate spectra - attachments 5-6 > per FEE64 p+n & nn 1.8.W spectra - attachments 7-8 > > 12.57 per FEE64 p+n & n+n 1.8.W spectra 20us FSR - attachments 9-10 > > Test signals > > EG&G Ortec 448 Pulser > Amplitude 90,000 > attenuation x10 > Normalise 5.00 > t_r 100ns > tau_d 50us > Frequency 50Hz ( relay line freq ) > Polarity + > > test + direct from 448, test - from Cooknell SA1 Sum & Invert > > Test + daisy chain -> aida01 - aida02 - aida05 - aida06 - 50 Ohm > Test - daisy chain -> aida03 - aida04 - aida07 - aida08 - 50 Ohm > > aida08 ( DSSSD #2 n+n Ohmic bias ) LK2 removed, LK1 added > aida04 ( DSSSD #1 n+n Ohmic bias ) no LK2 fitted, LK1 added > > DSSSD#1 SHV - 2x Lemo 00.250 cable replaced by standard SHV - BNC/BNC Lemo-00.250 coaxial cable > > Adaptor PCB ground daisy chain aida01 - aida02 - aida03 - aida04 > > DSSSD#2 - DSSSD#4 HV cabling unchanged > > 15.30 Power ON > System wide checks OK except > aida02 fail WR decoder status - attachment 11 > > FEE64 temperatures - attachment 12 > aida06 ASIC temp low > > DSSSD bias - attachment 13 > > ADC data item stats - attachment 14 > slow comparator 0x64 > > per FEE64 Rate and 1.8.W 20us FSR spectra - attachments 15-17 > > 16.00 per FEE64 rate spectra - attachment 18 > > 16.15 per FEE64 Rate and 1.8.W 20us FSR spectra - attachments 19-21 > > 17.36 Switch from VETAR2 ( all MACB modules mode 0x3 ) to NIM WR emulator ( root MACB module 0xd, all others 0x3 ) > All FEE64s fail Global clock and WR decoder - attachments 22-24 > WR timetsamp control - attachment 25 > > > 18.10 Noted MACB #4 mode actually set to 0x2 *not* 0x3 - changed to 0x3 > > Re-test with VETAR2 WR - same behaviour as observed this morning - see attachments 26-30 > > > > > Summary > > n+n sides appear to be OK but noisy > > no data observed from aida06 & aida11 > aida11 known to be disconnected from adaptor PCB > > aida15 asic #2 may be faulty > > aida01, aida05, aida12 - data suggest FEE64/adaptor PCB alignment issue ? > > aida02, aida03, aida09, aida14 - data suggest adaptor PCB damaged pins ? > > > aida07 - network config issue ? > > aida08 - not completing boot > > need to attach sys console to aida07 & aida08 to diagnose problem.
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown